# Simulation-based Approach for Fast Optimal Control of a Stefan Problem with Application to Cell Therapy

### v1

@@ A.1 Piecewise constant and linear controls
A.2 Number of control intervals
@@
The optimal control problems defined by (17) and (20) were solved via opt_Ipopt and $||e||_{2}$ was calculated for different values of $n_{c}$. Logically, the accuracy of the solution should improve, i.e., smaller $||e||_{2}$, when $n_{c}$ increases due to finer discretization. From Fig.  7, the accuracy of the solution does not change significantly after $n_{c}$ reaches some certain value, and so this threshold should be chosen for control vector parameterization. As a result, $n_{c}$ was set to 16 for Problem 2 (Fig.  7A) and 12 for Problem 3 (Fig.  7B).
@@
Appendix B Solver Options
There are a large number of options for optimization, ODE, and DAE solvers, which could lead to different accuracy and computational performance. This section describes and justifies the choice of important solver options used in this work. The default values were used for the options not mentioned here.
For all case studies, index-1 DAEs were solved by ode15s in MATLAB, whereas high-index DAEs were handled by GEKKO in Python. The differential index of DAEs was checked using reduceDAEIndex. The wall-clock times (aka wall times, clock times) were measured using the tic and toc functions, with each simulation repeated at least 10 times for consistent results, as indicated by the standard deviation of the measured wall times being smaller than 5%. For fair comparison of computational performance, the initial guess of $\Theta_{b}$ was set to the average value of 0.5 in all cases.
For the optimization-based approaches, the optimality tolerance was set to $10^{-7}$ for all optimization solvers, including IPOPT, fmincon, CasADi, and GEKKO. The integration tolerance for ode15s was set to $10^{-10}$. Using higher values for the optimality and integration tolerances could lead to inaccurate solutions, while using tighter values does not improve the accuracy significantly. For the objective functions (17) and (20) in Problems 2 and 3, the derivatives were approximated using a finite difference scheme, whereas the integrals can be approximated using a Riemann sum. The time step $\Delta\tau$ was set to 0.05.
While ode15s uses adaptive time-stepping, this option is not available in GEKKO; i.e., users have to specify a vector of time points for ODE integration. Hence, there is no ideal comparison between GEKKO’s ODE solvers and MATLAB’s ode15s. The fixed time step $\Delta\tau=0.05$ was specified for ODE integration in GEKKO, with five collocation points (INODES = 5) for each time interval. This time step was chosen to be consistent with the value used for a finite difference and Riemann sum approximation mentioned in the previous paragraph, while the number of collocation points was selected such that that the temperature and interface position simulated by GEKKO’s ODE solver have the same level of accuracy as those simulated by ode15s.
For the simulation-based approach, ode15s was used for an index-1 system (Problem 1), with the exact same solver options as mentioned above. GEKKO’s DAE solver was used for high-index systems (Problems 2 and 3). As a reformulated DAE system is different from the original ODE system, the time step needs to be adjusted accordingly. The time step was selected such that the accuracy of the simulation-based solution measured by $||e||_{2}$ is on the same order of magnitude as that of the optimization-based solution, giving $\Delta\tau=0.12$ and $\Delta\tau=0.37$ for Problems 2 and 3, respectively.
The accuracy of both the optimization- and simulation-based approaches is measured using $||e||_{2}$ defined by (19) and (22). To calculate $||e||_{2}$, the derivatives were approximated using a finite difference scheme with $\Delta\tau=0.05$.
Appendix C Bounds on the Interface Velocity and Rate of Change in Temperature
We show in Fig.  2, Problems 2 and 3, and Section 4.4 that, if the target velocity (and also the rate of temperature change) is chosen properly, the bound on a control vector will never be violated. A formal proof is given in this section.
To obtain a closed-form solution for the proof, the heat conduction in the liquid domain is assumed to be quasi-steady. This approximation is accurate for problems associated with phase change as most of the heat transfer occurs at the moving interface. In this case, the closed-form solution for the interface velocity is
@@ Appendix C Bounds on the Interface Velocity and Rate of Change in Temperature
where all the parameter description and values can be found in Srisuma et al. (2023b). From (23), the interface velocity $(dS/d\tau)$ is a monotonic function of the heater temperature $(\Theta_{b})$. Hence, the condition in which the bounds on the heater temperature are not violated can be obtained by substituting the upper bound $\Theta_{b}=1$ into (23), which is
$\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}}>-\frac{k_{2}U(T_{0}-T_{m})}{\rho\alpha_{1}\Delta H_{\!f}S\left(U\ln(1/S)+k/b\right)}.$ (24)
With the parameter values in Srisuma et al. (2023b), the right-hand side of (24) is about $-0.13$. This conclusion agrees with the results presented in Problems 3 and 4, in which the bounds are not violated when $(dS/d\tau)_{\mathrm{sp}}$ is higher than $-0.13$, whereas the upper bound is active when $(dS/d\tau)_{\mathrm{sp}}$ is $-0.15$. A similar strategy can be used for analyzing the temperature change, e.g., using a thermal lumped capacity approximation.

### v2

@@ B.1 Piecewise constant and linear controls
B.2 Number of control intervals
@@
The optimal control problems defined by (17) and (20) are solved via opt_Ipopt and $||e||_{2}$ is calculated for different values of $n_{c}$. Logically, the accuracy of the solution should improve (i.e., $||e||_{2}$ decreases) when $n_{c}$ increases, due to finer discretization. From Fig. 7, the accuracy of the solution does not change significantly after $n_{c}$ reaches some certain value, and so this threshold should be chosen for control vector parameterization. As a result, $n_{c}$ is set to 16 for Problem 2 (Fig.  7A) and 12 for Problem 3 (Fig. 7B).
@@
Appendix C Bounds on the Interface Velocity and Rate of Temperature Change
We show in Fig. 2, Problems 2 and 3, and Section 4.4 that, if the target velocity (and also the rate of temperature change) is chosen properly, the bound on a control vector will never be violated. A formal proof is given in this section.
To obtain a closed-form solution for the proof, heat conduction in the liquid domain is assumed to be quasi-steady. This approximation is accurate for problems associated with phase change in general as the sensible heat is much smaller than the latent heat. In this case, the closed-form solution for the interface velocity is
@@ Appendix C Bounds on the Interface Velocity and Rate of Temperature Change
where the parameter values and description can be found in Srisuma et al. (2023b). From (23), the interface velocity $dS/d\tau$ is a monotonic function of the heater temperature $\Theta_{b}$. Substituting the upper bound $\Theta_{b}=1$ into (23) yields
$\left(\dfrac{dS}{d\tau}\right)_{\!\!\mathrm{sp}}>-\frac{k_{2}U(T_{0}-T_{m})}{\rho\alpha_{1}\Delta H_{\!f}S\left(U\ln(1/S)+k/b\right)},$ (24)
which gives the limit on the target velocity that can be specified without violating the upper bound on the heater temperature. With the parameter values in Srisuma et al. (2023b), the right-hand side of (24) is about $-0.13$. This conclusion agrees with the results presented in Problems 3 and 4, in which the bounds are not violated when $(dS/d\tau)_{\mathrm{sp}}$ is higher than $-0.13$, whereas the upper bound is active when $(dS/d\tau)_{\mathrm{sp}}$ is $-0.15$. A similar strategy can be used to analyze the temperature change, e.g., using a thermal lumped capacity approximation.
@@
Data Availability
Software and data used in this work are available at https://github.com/PrakitrSrisuma/simDAE-optimalcontrol.